vidyo

vidyo link PIN 1055

Action Items:

News:

Franco: Application for the DT upgrade for phase 2 submitted.

Tommaso: INFN will open ~70 positions, 15 will be to reserved to technological positions (tecnologi).

Tommaso: this Thursday 10am Benjamin Annach, well known theorist will give a talk in "Aula A".

RobertoR: 100 students back from CERN. We have been able to visit Compass, Alice and CMS. Visit successful, students were enthusiast, we hope to see some of them in a year from now asking for a thesis. Applied for G&V funds for Luca Bonaldo, a PD student who just graduated as a string theorist but who wanted to come to work with us.

Framework & production

Federica: no news about framework. Deleting datasets. 318TB free in Legnaro. We have a tremendous amount of space now.

On Thursday or Friday there was a peak in the user interface, interactive process producing a lot of I/O.

Analysis

B physics

Stefano: sent to PhysCoord and spokesperson the proposal for open trigger for unbiased samples. Reception good, but more studies needed. Stefano can circulate the paper. Triggering 1 mu on 1 side, getting anything on the other (bbbar production). This will be parked data, the event might be fully stored, or only part of the FEDs can be saved, still to be studied/discussed.

Still discussing with Mauro the answer to the comments to the PLB paper to be published.

Alberto: Wrote the code for the soft muon MVA ID in CMSSW, is been imported in the ID code. It will be available as a flag to active. This might be work used to get ~1month worth EPR.

Charged Higgs to top-beauty

Jacopo: replied to requests raised during pre-approval. One last Q about expected limits with different approaches. Will reply tomorrow or the day after.

Dibosons, VZ resonance (V=W,Z)

Jacopo: leptonic analysis 2l2q, submitted the aswers to CWR 2 weeks ago. Pin

Invisble +: Final reading

4B full hadronic

Siew: Still doubts from the conveners about the BSM points. One possibility is to use ONLY the SM trained BDT, NOT the ones trained on the BSM ponts.

They are asking to quantify disagreements observed during some chackes done in the past. Discussing with them on the possibility to measure the disagreement and maybe turn the disagreement into an increase in systematics.

Higgs Dark Matter

2 news. There is a strong cutoff on the 2016 data, there is a strong motivation from our side, we use a 1.5 cone size, we have a double b-tagger calibrated ONLY on 2016 data. The conveners ask for a check, to run the anaysis w/o the double b-tagger. check has been done and circulated inside the working group before sending the results to the conveners.

Jacopo and Marco working on spark.

EPR

Franco: please subscribe

Trigger(less)

Jacopo: at CERN we worked with Triossi, we sent data from the PCI express to the memory of the PC using the firmware Andrea wrote. Next step is to thest the whole chain, from the DT to the PC.

Detector

Luminosity & Bril

Reinstallation on the - side of the detector. Some of the fibres are a little dirty.

Analysis: working.

-- Main.Franco Simonetto 04-10-2017

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2018-02-26 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback